# Structural Prognostic Event Modeling for Multimodal Cancer Survival Analysis

### v1

@@ 4.1 Datasets and Implementations
We perform comprehensive evaluations across ten publicly available cancer cohorts from TCGA111https://portal.gdc.cancer.gov/. Case counts for each dataset are listed in Table 1. Following (jaume2024modeling), we predict disease-specific survival (DSS) with histological data including all diagnostic WSIs and transcriptomic profiles with DSS labels are obtained from cBioPortal222https://www.cbioportal.org/. Pathway gene sets are curated from Hallmarks (subramanian2005gene; liberzon2015molecular) and Reactome (gillespie2022reactome), with genes absent in cBioPortal removed. We use 5-fold cross-validation and report the concordance index (C-index) (harrell1996multivariable) with standard deviation (std) and visualize the Kaplan–Meier (KM) kaplan1958nonparametric survival curves for stratified risk groups, perform the log-rank test mantel1966evaluation and restricted mean survival time (RMST) (irwin1949standard; karrison1986restricted) to assess differences between them. Further information can be found in Appendix C.
@@ 4.2 Comparisons with SOTAs
We benchmark our approach against three categories of state-of-the-art (SOTA) methods: (1) Transcriptomics. We consider a feed-forward neural network (MLP) (haykin1998neural), a pathway-specific baseline (SNNs) (klambauer2017self), and SNNTrans, which integrates SNNs with transformers (klambauer2017self; shao2021transmil). (2) Histology. We compare with leading MIL-based models, including ABMIL (ilse2018attention), TransMIL (shao2021transmil), and CLAM (lu2021data). (3) Multimodal. We evaluate against eight multimodal SOTA methods including Porpoise (chen2022pan), MCAT (chen2021multimodal), MOTCat (xu2023multimodal), CMTA (zhou2023cross), SurvPath (jaume2024modeling), PIBD (zhang2024prototypical), MMP song2024multimodal, and LD-CVAE (zhou2025robust). Among them, the first five methods adopt instance-level interactions for histology and omics integration, PIBD and MMP employ prototype-based modeling, while LD-CVAE is specifically designed to handle missing genomic data.
@@ 4.3 Robustness
Since collecting genomic data is substantially more costly than acquiring histology images (payne2018cost), clinical deployment often faces performance degradation when genomic features are unavailable. To evaluate the robustness of our approach, we test SlotSPE under the missing-genomic-modality setting (Table 2). SlotSPE maintains strong performance, and in several cohorts (e.g., HNSC, STAD) its missing-modality variant matches or even surpasses the full-modality version, demonstrating resilience to incomplete omics data. This robustness stems from the cross-modal reconstruction task, which enables the model to infer pathway representations from histology through omics-derived slots. Compared with LD-CVAE (zhou2025robust)—a method explicitly designed for missing genomic data—SlotSPE shows a somewhat larger drop relative to its full-modality variant, yet still achieves superior or comparable performance across most datasets. Notably, its overall performance even exceeds that of full-modality LD-CVAE. These findings highlight that SlotSPE not only integrates multimodal features effectively but also delivers robust predictions in the absence of genomic data.
@@ 4.6 Efficiency
In this section, we assess the inference efficiency of our model by comparing it against the four strongest baselines listed in Table 1. Because all methods share the same WSI and omics encoders, we exclude these components and evaluate only the model-specific modules to ensure a fair and meaningful comparison. We measure peak inference memory and wall-clock time on the COADREAD dataset, using all WSI patches during inference and configuring SlotSPE with 16 slots and 3 iterations. As shown in Figure 4, SlotSPE achieves the highest C-index while maintaining low runtime and memory usage. MCAT and MOTCAT achieve shorter runtime primarily because they omit self-attention computations within WSI patches; however, this design choice leads to a considerable drop in predictive performance. CMTA improves upon these baselines by employing Nyström attention (xiong2021nystromformer) as an approximation to full self-attention, but its capacity remains insufficient to capture the sparse, patient-specific structural prognostic signals. LD-CVAE attains the second-best performance, yet its reliance on a variational autoencoder introduces substantial memory and runtime overhead. Overall, SlotSPE offers the most favorable balance between efficiency and predictive performance, with minimal memory usage and the second-fastest runtime among all compared methods. Additional analysis is provided in the Appendix G.
@@ 4.7 Interpretability
Cross-modality Alignment. Visualization of slot assignments (Figure 5A) shows that histology- and genomics-derived slots often align to structurally similar tissue regions in WSIs, reflecting event-level cross-modal consistency. Subtle differences remain, which can be attributed to modality-specific information: morphology-driven slots emphasize structural patterns, whereas genomics-driven slots highlight regions linked to molecular features. Intra- and Inter-modal Interactions. For intra-modal analysis, we visualize attention maps of histology-derived slots (first three plots in Figure 5B) and report the top-3 most relevant and irrelevant pathways per omics-derived slot (Figure 5C). For inter-modal analysis, we link pathway-level attention scores with the corresponding genomic-slot attention maps on WSIs (last three plots in Figure 5B). For instance, in Slot 2, highly attended pathways correspond to tumor-enriched regions, consistent with prior studies on breast cancer prognosis (wang2010eicosanoids; neman2014human; chen2019loss; ta2021novel; xie2023systematic).
In general, these interpretability analyzes demonstrate the ability of SlotSPE to establish event-level alignment across modalities and reveal pathway–morphology correspondences, offering the potential for biologically plausible explanations and new hypotheses for further validation. Additional examples are provided in the Appendix I.
@@ B.1 Problem Formulation
Given the -th patient with multimodal inputs where denotes the pathology data, denotes the genomic profile, is the discretized survival time interval (risk band), and is the censoring indicator ( for an observed event, for right-censoring). Following the multiple instance learning (MIL) paradigm (chen2021multimodal; jaume2024modeling), features extracted from the WSI patch encoder and gene encoder are organized as bags of instances: and , where and denote the numbers of WSI patches and biological pathways, respectively, and is the feature dimension. The objective is to learn mapping for survival prediction. We denote by the predicted hazard probability at a discrete time interval , i.e., the probability of an event occurring at given survival up to . The survival function is . With patients in the training set, the hazard, survival, and the negative log-likelihood (NLL) loss (zadeh2020bias) are:
@@ B.2 Details of Selective Slot Activation
To achieve sparse yet differentiable slot selection, we adopt the Gumbel-Top- (gumbel1954statistical; maddison2014sampling; kool2019stochastic) trick combined with a Straight-Through (ST) estimator (jang2016categorical).
@@ B.3 Details of Reconstructions
Slots Regularization as Reconstruction. Within each modality, slots are required to reconstruct the original modality-specific features to address the emergence of null slots. We use a Transformer decoder (vaswani2017attention) as the reconstruction head. For genomics, slot embeddings are decoded to approximate the original pathway embeddings , guided by the positional embeddings (Eq. 5), with reconstruction optimized via MSE loss:
@@ B.4 Details of Multimodal Interactions
In contrast to one-shot co-attention methods (chen2021multimodal; xu2023multimodal; jaume2024modeling), we employ an iterative cross-attention scheme that reuses a single fusion layer for iterations of bidirectional updates. At each iteration , omics- and histology-derived slots alternately attend to one another, absorbing complementary information while preserving their own context via recurrent updates, akin to slot attention. This iterative design supports multi-step fusion, progressively aligning molecular and morphological evidence that may require hierarchical inference (e.g., molecular dysregulation propagating through cellular morphology to tissue-level structure (huang2025bridging)). Concretely, at iteration , the update of histology slots attending to omics slots is defined as
@@ Appendix C Datasets and Implementations
Dataset. We perform comprehensive evaluations across ten publicly available cancer cohorts from TCGA333https://portal.gdc.cancer.gov/ , including Breast Invasive Carcinoma (BRCA), Bladder Urothelial Carcinoma (BLCA), Colon and Rectum Adenocarcinoma (COADREAD), Stomach Adenocarcinoma (STAD), Head and Neck Squamous Cell Carcinoma (HNSC), as well as Lung Adenocarcinoma (LUAD), Lung Squamous Cell Carcinoma (LUSC), Kidney Renal Clear Cell Carcinoma (KIRC), and Skin Cutaneous Melanoma (SKCM). Case counts for each dataset are listed in Table 1. Following (jaume2024modeling), we predict disease-specific survival (DSS), a more precise indicator of patient status than overall survival. Histological data include all diagnostic WSIs, while transcriptomic profiles with DSS labels are obtained from cBioPortal444https://www.cbioportal.org/. Pathway gene sets are curated from Hallmarks (subramanian2005gene; liberzon2015molecular) and Reactome (gillespie2022reactome), with genes absent in cBioPortal removed, yielding 330 pathways.
Evaluation. We adopt 5-fold cross-validation and report the concordance index (C-index) (harrell1996multivariable) together with the standard deviation (std) across folds. To further assess clinical relevance, we visualize Kaplan–Meier (KM) survival curves (kaplan1958nonparametric) for stratified risk groups and apply the log-rank test (mantel1966evaluation) to evaluate global differences in survival distributions. In addition, we compute the restricted mean survival time (RMST) (irwin1949standard; karrison1986restricted). RMST, defined as the area under the estimated survival curve up to a clinically meaningful truncation time (60 months in our experiments), provides an interpretable summary of the average survival time within the specified horizon and does not rely on the proportional hazards assumption. We report: (a) the RMST values for both high- and low-risk groups, (b) the RMST difference ( High–Low) with a 95% bootstrap confidence interval (CI) and a two-sided -value testing whether the difference is significantly different from zero, and (c) the RMST ratio (High/Low) with a log-transformed 95% bootstrap CI as a scale-free measure of group separation. To ensure robust inference, both CI and -values are estimated from 1000 bootstrap replicates (tibshirani1993introduction). These statistics are annotated on the KM plots, providing an interpretable and robust evaluation of prognostic stratification.
@@
Implementation. Tissue regions are segmented from WSIs, and non-overlapping patches are extracted at magnification. We adopt UNI (chen2024towards) as the patch encoder and SNNs as the pathway encoder. The method is implemented in Python using PyTorch and executed on only one NVIDIA A100 GPU. For survival prediction, disease-specific survival (DSS) time is discretized into four intervals. Bag features are embedded via an MLP with a 512-dimensional hidden layer and a 256-dimensional output, thus and are all set to 256. The number of iterations for slot attention is set to , and for iterative cross-attention to . The top- in selective slot activation decoder is set to slot numbers since patients are divided into 4 risk groups. We report the best performance by ablating the slot numbers in Appendix F.3.1. Pathway tokens are reconstructed using MSE loss, while histopathology tokens are optimized with cosine similarity loss using a hyperparameter (searched from ). To support batch training, 4096 patches are randomly sampled per WSI, whereas all patches are used during inference. Optimization is carried out with Adam (kingma2014adam) at a learning rate of . All models are trained for 30 epochs with a batch size of 32.
@@ Appendix D Comparisons with SOTAs
To further assess the generalization ability of our approach, we compare SlotSPE with other multimodal methods under a higher-level grouping of genes into six functional categories (subramanian2005gene; liberzon2015molecular; chen2021multimodal; xu2023multimodal; zhou2023cross): tumor supression, oncogenesis, protein kinases, cellular differentiation, transcription, and cytokines and growth. In this setting, the number of gene slots is fixed to six, while the number of histopathology slots is varied and the best performance is reported. Results in Table 5 show that SlotSPE achieves the best overall performance, ranking first or second in 8 out of 10 cohorts, and outperforming the second-best method (LD-CVAE) by 1.9% on average. This indicates that even when genes are grouped into higher-level functional categories, our SlotSPE is able to maintain strong predictive performance across different granularities of gene input.
@@ Appendix E Robustness and Risk Stratification
The complete results with standard deviations for Table 2, using the biological pathway representation for gene data, are reported in Table 6. While most baselines are not explicitly designed to handle missing modalities, they can still be evaluated under missing-genomics settings by supplying a neutral placeholder input in place of genomic features (zhang2025adamhf). For completeness, we also include two strong methods (MOTCAT and CMTA). The results show that when these approaches rely solely on histology—without modeling or reconstructing genomic information—their prognostic performance drops substantially. Moreover, under missing-genomics conditions, their limited robustness leads to performance even worse than strong single-modal baselines (e.g., ABMIL).
@@ F.2 Histology Encoder ablations
To assess the robustness of SlotSPE to the choice of visual encoder, we first replace the histology encoder with a ResNet50 (srivastava2015highway; he2016deep) pretrained only on ImageNet (imagenet_cvpr09), rather than a pathology-specific foundation model. This setting tests whether SlotSPE can still achieve strong performance when the visual backbone lacks domain-specific pretraining. For this experiment, we evaluate representative unimodal and multimodal baselines listed in Table 1. As summarized in Table 9, SlotSPE continues to outperform all baselines under this weaker configuration, suggesting that its performance is not tightly coupled to a specialized encoder. We further extend this analysis by examining SlotSPE with recent pathology foundation models, including CONCH (lu2024visual), CONCH v1.5 (used in TITAN (ding2025multimodal)), and UNI (chen2024towards). The results in Figure 7 show that performance varies across different encoders: while recent pathology foundation models generally improve results compared to ImageNet-pretrained encoder, the best performance is obtained with UNI. These results indicate that SlotSPE is both robust to weaker, non-specialized encoders and compatible with strong pathology-specific foundation models.

### v2

@@ 4.1 Datasets and Implementations
We perform comprehensive evaluations across ten publicly available cancer cohorts from TCGA111https://portal.gdc.cancer.gov/. Case counts for each dataset are listed in Table 1. Following (Jaume et al., 2024), we predict disease-specific survival (DSS) with histological data including all diagnostic WSIs and transcriptomic profiles with DSS labels are obtained from cBioPortal222https://www.cbioportal.org/. Pathway gene sets are curated from Hallmarks (Subramanian et al., 2005; Liberzon et al., 2015) and Reactome (Gillespie et al., 2022), with genes absent in cBioPortal removed. We use 5-fold cross-validation and report the concordance index (C-index) (Harrell Jr et al., 1996) with standard deviation (std) and visualize the Kaplan–Meier (KM) Kaplan and Meier (1958) survival curves for stratified risk groups, perform the log-rank test Mantel and others (1966) and restricted mean survival time (RMST) (Irwin, 1949; Karrison, 1986) to assess differences between them. Further information can be found in Appendix C.
@@ 4.2 Comparisons with SOTAs
We benchmark our approach against three categories of state-of-the-art (SOTA) methods: (1) Transcriptomics. We consider a feed-forward neural network (MLP) (Haykin, 1998), a pathway-specific baseline (SNNs) (Klambauer et al., 2017), and SNNTrans, which integrates SNNs with transformers (Klambauer et al., 2017; Shao et al., 2021). (2) Histology. We compare with leading MIL-based models, including ABMIL (Ilse et al., 2018), TransMIL (Shao et al., 2021), and CLAM (Lu et al., 2021). (3) Multimodal. We evaluate against eight multimodal SOTA methods including Porpoise (Chen et al., 2022), MCAT (Chen et al., 2021b), MOTCat (Xu and Chen, 2023), CMTA (Zhou and Chen, 2023), SurvPath (Jaume et al., 2024), PIBD (Zhang et al., 2024), MMP Song et al. (2024), and LD-CVAE (Zhou et al., 2025). Among them, the first five methods adopt instance-level interactions for histology and omics integration, PIBD and MMP employ prototype-based modeling, while LD-CVAE is specifically designed to handle missing genomic data.
@@ 4.3 Robustness
Since collecting genomic data is substantially more costly than acquiring histology images (Payne et al., 2018), clinical deployment often faces performance degradation when genomic features are unavailable. To evaluate the robustness of our approach, we test SlotSPE under the missing-genomic-modality setting (Table 2). SlotSPE maintains strong performance, and in several cohorts (e.g., HNSC, STAD) its missing-modality variant matches or even surpasses the full-modality version, demonstrating resilience to incomplete omics data. This robustness stems from the cross-modal reconstruction task, which enables the model to infer pathway representations from histology through omics-derived slots. Compared with LD-CVAE (Zhou et al., 2025)—a method explicitly designed for missing genomic data—SlotSPE shows a somewhat larger drop relative to its full-modality variant, yet still achieves superior or comparable performance across most datasets. Notably, its overall performance even exceeds that of full-modality LD-CVAE. These findings highlight that SlotSPE not only integrates multimodal features effectively but also delivers robust predictions in the absence of genomic data.
@@ 4.6 Efficiency
In this section, we assess the inference efficiency of our model by comparing it against the four strongest baselines listed in Table 1. Because all methods share the same WSI and omics encoders, we exclude these components and evaluate only the model-specific modules to ensure a fair and meaningful comparison. We measure peak inference memory and wall-clock time on the COADREAD dataset, using all WSI patches during inference and configuring SlotSPE with 16 slots and 3 iterations. As shown in Figure 4, SlotSPE achieves the highest C-index while maintaining low runtime and memory usage. MCAT and MOTCAT achieve shorter runtime primarily because they omit self-attention computations within WSI patches; however, this design choice leads to a considerable drop in predictive performance. CMTA improves upon these baselines by employing Nyström attention (Xiong et al., 2021) as an approximation to full self-attention, but its capacity remains insufficient to capture the sparse, patient-specific structural prognostic signals. LD-CVAE attains the second-best performance, yet its reliance on a variational autoencoder introduces substantial memory and runtime overhead. Overall, SlotSPE offers the most favorable balance between efficiency and predictive performance, with minimal memory usage and the second-fastest runtime among all compared methods. Additional analysis is provided in the Appendix G.
@@ 4.7 Interpretability
Cross-modality Alignment. Visualization of slot assignments (Figure 5A) shows that histology- and genomics-derived slots often align to structurally similar tissue regions in WSIs, reflecting event-level cross-modal consistency. Subtle differences remain, which can be attributed to modality-specific information: morphology-driven slots emphasize structural patterns, whereas genomics-driven slots highlight regions linked to molecular features. Intra- and Inter-modal Interactions. For intra-modal analysis, we visualize attention maps of histology-derived slots (first three plots in Figure 5B) and report the top-3 most relevant and irrelevant pathways per omics-derived slot (Figure 5C). For inter-modal analysis, we link pathway-level attention scores with the corresponding genomic-slot attention maps on WSIs (last three plots in Figure 5B). For instance, in Slot 2, highly attended pathways correspond to tumor-enriched regions, consistent with prior studies on breast cancer prognosis (Wang and DuBois, 2010; Neman et al., 2014; Chen et al., 2019; Ta et al., 2021; Xie et al., 2023). In general, these interpretability analyzes demonstrate the ability of SlotSPE to establish event-level alignment across modalities and reveal pathway–morphology correspondences, offering the potential for biologically plausible explanations and new hypotheses for further validation. Additional examples are provided in the Appendix I.
@@ B.1 Problem Formulation
Given the -th patient with multimodal inputs where denotes the pathology data, denotes the genomic profile, is the discretized survival time interval (risk band), and is the censoring indicator ( for an observed event, for right-censoring). Following the multiple instance learning (MIL) paradigm (Chen et al., 2021b; Jaume et al., 2024), features extracted from the WSI patch encoder and gene encoder are organized as bags of instances: and , where and denote the numbers of WSI patches and biological pathways, respectively, and is the feature dimension. The objective is to learn mapping for survival prediction. We denote by the predicted hazard probability at a discrete time interval , i.e., the probability of an event occurring at given survival up to . The survival function is . With patients in the training set, the hazard, survival, and the negative log-likelihood (NLL) loss (Zadeh and Schmid, 2020) are:
@@ B.2 Details of Selective Slot Activation
To achieve sparse yet differentiable slot selection, we adopt the Gumbel-Top- (Gumbel, 1954; Maddison et al., 2014; Kool et al., 2019) trick combined with a Straight-Through (ST) estimator (Jang et al., 2016).
@@ B.3 Details of Reconstructions
Slots Regularization as Reconstruction. Within each modality, slots are required to reconstruct the original modality-specific features to address the emergence of null slots. We use a Transformer decoder (Vaswani et al., 2017) as the reconstruction head. For genomics, slot embeddings are decoded to approximate the original pathway embeddings , guided by the positional embeddings (Eq. 5), with reconstruction optimized via MSE loss:
@@ B.4 Details of Multimodal Interactions
In contrast to one-shot co-attention methods (Chen et al., 2021b; Xu and Chen, 2023; Jaume et al., 2024), we employ an iterative cross-attention scheme that reuses a single fusion layer for iterations of bidirectional updates. At each iteration , omics- and histology-derived slots alternately attend to one another, absorbing complementary information while preserving their own context via recurrent updates, akin to slot attention. This iterative design supports multi-step fusion, progressively aligning molecular and morphological evidence that may require hierarchical inference (e.g., molecular dysregulation propagating through cellular morphology to tissue-level structure (Huang et al., 2025)). Concretely, at iteration , the update of histology slots attending to omics slots is defined as
@@ Appendix C Datasets and Implementations
Dataset. We perform comprehensive evaluations across ten publicly available cancer cohorts from TCGA333https://portal.gdc.cancer.gov/ , including Breast Invasive Carcinoma (BRCA), Bladder Urothelial Carcinoma (BLCA), Colon and Rectum Adenocarcinoma (COADREAD), Stomach Adenocarcinoma (STAD), Head and Neck Squamous Cell Carcinoma (HNSC), as well as Lung Adenocarcinoma (LUAD), Lung Squamous Cell Carcinoma (LUSC), Kidney Renal Clear Cell Carcinoma (KIRC), and Skin Cutaneous Melanoma (SKCM). Case counts for each dataset are listed in Table 1. Following (Jaume et al., 2024), we predict disease-specific survival (DSS), a more precise indicator of patient status than overall survival. Histological data include all diagnostic WSIs, while transcriptomic profiles with DSS labels are obtained from cBioPortal444https://www.cbioportal.org/. Pathway gene sets are curated from Hallmarks (Subramanian et al., 2005; Liberzon et al., 2015) and Reactome (Gillespie et al., 2022), with genes absent in cBioPortal removed, yielding 330 pathways.
Evaluation. We adopt 5-fold cross-validation and report the concordance index (C-index) (Harrell Jr et al., 1996) together with the standard deviation (std) across folds. To further assess clinical relevance, we visualize Kaplan–Meier (KM) survival curves (Kaplan and Meier, 1958) for stratified risk groups and apply the log-rank test (Mantel and others, 1966) to evaluate global differences in survival distributions. In addition, we compute the restricted mean survival time (RMST) (Irwin, 1949; Karrison, 1986). RMST, defined as the area under the estimated survival curve up to a clinically meaningful truncation time (60 months in our experiments), provides an interpretable summary of the average survival time within the specified horizon and does not rely on the proportional hazards assumption. We report: (a) the RMST values for both high- and low-risk groups, (b) the RMST difference ( High–Low) with a 95% bootstrap confidence interval (CI) and a two-sided -value testing whether the difference is significantly different from zero, and (c) the RMST ratio (High/Low) with a log-transformed 95% bootstrap CI as a scale-free measure of group separation. To ensure robust inference, both CI and -values are estimated from 1000 bootstrap replicates (Tibshirani and Efron, 1993). These statistics are annotated on the KM plots, providing an interpretable and robust evaluation of prognostic stratification.
@@
Implementation. Tissue regions are segmented from WSIs, and non-overlapping patches are extracted at magnification. We adopt UNI (Chen et al., 2024) as the patch encoder and SNNs as the pathway encoder. The method is implemented in Python using PyTorch and executed on only one NVIDIA A100 GPU. For survival prediction, disease-specific survival (DSS) time is discretized into four intervals. Bag features are embedded via an MLP with a 512-dimensional hidden layer and a 256-dimensional output, thus and are all set to 256. The number of iterations for slot attention is set to , and for iterative cross-attention to . The top- in selective slot activation decoder is set to slot numbers since patients are divided into 4 risk groups. We report the best performance by ablating the slot numbers in Appendix F.3.1. Pathway tokens are reconstructed using MSE loss, while histopathology tokens are optimized with cosine similarity loss using a hyperparameter (searched from ). To support batch training, 4096 patches are randomly sampled per WSI, whereas all patches are used during inference. Optimization is carried out with Adam (Kingma and Ba, 2014) at a learning rate of . All models are trained for 30 epochs with a batch size of 32.
@@ Appendix D Comparisons with SOTAs
To further assess the generalization ability of our approach, we compare SlotSPE with other multimodal methods under a higher-level grouping of genes into six functional categories (Subramanian et al., 2005; Liberzon et al., 2015; Chen et al., 2021b; Xu and Chen, 2023; Zhou and Chen, 2023): tumor supression, oncogenesis, protein kinases, cellular differentiation, transcription, and cytokines and growth. In this setting, the number of gene slots is fixed to six, while the number of histopathology slots is varied and the best performance is reported. Results in Table 5 show that SlotSPE achieves the best overall performance, ranking first or second in 8 out of 10 cohorts, and outperforming the second-best method (LD-CVAE) by 1.9% on average. This indicates that even when genes are grouped into higher-level functional categories, our SlotSPE is able to maintain strong predictive performance across different granularities of gene input.
@@ Appendix E Robustness and Risk Stratification
The complete results with standard deviations for Table 2, using the biological pathway representation for gene data, are reported in Table 6. While most baselines are not explicitly designed to handle missing modalities, they can still be evaluated under missing-genomics settings by supplying a neutral placeholder input in place of genomic features (Zhang et al., 2025). For completeness, we also include two strong methods (MOTCAT and CMTA). The results show that when these approaches rely solely on histology—without modeling or reconstructing genomic information—their prognostic performance drops substantially. Moreover, under missing-genomics conditions, their limited robustness leads to performance even worse than strong single-modal baselines (e.g., ABMIL).
@@ F.2 Histology Encoder ablations
To assess the robustness of SlotSPE to the choice of visual encoder, we first replace the histology encoder with a ResNet50 (Srivastava et al., 2015; He et al., 2016) pretrained only on ImageNet (Deng et al., 2009), rather than a pathology-specific foundation model. This setting tests whether SlotSPE can still achieve strong performance when the visual backbone lacks domain-specific pretraining. For this experiment, we evaluate representative unimodal and multimodal baselines listed in Table 1. As summarized in Table 9, SlotSPE continues to outperform all baselines under this weaker configuration, suggesting that its performance is not tightly coupled to a specialized encoder. We further extend this analysis by examining SlotSPE with recent pathology foundation models, including CONCH (Lu et al., 2024), CONCH v1.5 (used in TITAN (Ding et al., 2025)), and UNI (Chen et al., 2024). The results in Figure 7 show that performance varies across different encoders: while recent pathology foundation models generally improve results compared to ImageNet-pretrained encoder, the best performance is obtained with UNI. These results indicate that SlotSPE is both robust to weaker, non-specialized encoders and compatible with strong pathology-specific foundation models.